ASTROCHEMISTRY WITH THE UMIST DATABASE 2005

J. WOODALL, Jodrell Bank Centre for Astrophysics, School of Physics and Astronomy, University of Manchester, Sackville Street Building, Manchester M60 1QD, UK; M. AGÚNDEZ, Dpto. Astrofisica Molecular e Infrarroja, Instituto de Estructura de la Materia (CSIC), C Serrano 121 28806, Madrid, Spain; \underline{A. J. MARKWICK-KEMPER}, Department of Astronomy, University of Virginia, Charlottesville, VA 22904, USA and National Radio Astronomy Observatory, 520 Edgemont Road, Charlottesville, VA 22903, USA; T. J. MILLAR, Astrophysics Research Centre, School of Mathematics and Physics, Queen’s University Belfast, Belfast BT7 1NN, UK.

A new version of the UMIST Database for Astrochemistry is available, the fourth such version to be released to the community. The UDfA 2005 contains 4566 binary gas-phase reactions, an increase of 10% from the previous (1999) version. Major updates have been made to ion-neutral reactions, neutral-neutral reactions (particularly at low temperature), and dissociative recombination reactions. We have also augmented the chemistry to include fluorine. The database, associated software and documentation are to be found online at www.udfa.net.